### v1

@@ Abstract
We propose a model for the coupling of flow and transport equations with porous membrane-type conditions on part of the boundary. The governing equations consist of the incompressible Navier–Stokes equations coupled with an advection-diffusion equation, and we employ a Lagrange multiplier to enforce the coupling between penetration velocity and transport on the membrane, while mixed boundary conditions are considered in the remainder of the boundary. We show existence and uniqueness of the continuous problem using a fixed-point argument. Next, an H(div)-conforming finite element formulation is proposed, and we address its a priori error analysis. The method uses an upwind approach that provides stability in the convection-dominated regime. We showcase a set of numerical examples validating the theory and illustrating the use of the new methods in the simulation of reverse osmosis processes.
@@
Navier–Stokes equations coupled with transport; Lagrange multipliers; Reverse osmosis; Divergence-conforming finite element methods.
@@ 1 Introduction
1.1 Scope
@@
The coupling of Navier–Stokes equations and advection-diffusion equations is central in many applications in industry and engineering. One of such instances is the simulation of filtration processes that occur in water purification where high velocity flow of water with relatively high concentration of salt goes through a unit under reverse osmosis. Even without considering the mechanisms of ion transport and molecule interactions through the membrane, the sole fluid dynamics process already poses interesting open questions. In that context, we can consider a simplified model where water and salt transport are coupled through advection and through a boundary condition at the membrane stating that the filtration velocity is proportional to a linear function of the salt concentration. This model has been recently studied numerically in [10], where the coupling mechanisms on the membrane are incorporated through a Nitsche approach. Here we rewrite that model using a membrane Lagrange multiplier and provide a complete well-posedness analysis as well as the analysis of a pressure-robust discretisation. In analysing the coupled system and separating for a moment the salt transport from the incompressible flow equations, we end up with a generalisation of the Navier–Stokes equations with slip boundary conditions which have been studied extensively starting from the conforming finite element methods proposed in [41, 42]. The off-diagonal bilinear form is different from the usual one in that it also includes a pairing between the normal trace of velocity and the Lagrange multiplier taking into account the membrane coupling. Another difficulty in the model considered herein is that a non-homogeneous boundary condition is needed at the inlet. Often the analysis is simply restricted to the case of homogeneous essential boundary conditions, but here the non-homogeneity is important as it permits that the coupling occurs (otherwise the membrane coupling vanishes, and the only solution is the trivial one). We note that for smooth domains it is expected that discretisations are prone to see the so-called Babuška paradox – a variational crime associated with the approximation of the boundary, and where a sub-optimal convergence is expected (see more details in, e.g., [26, 40]). Similar works focusing on slip boundary conditions imposed with Lagrange multipliers or with penalty can be found in [32, 30, 43]. However in our case we focus the analysis to the case of polygonal boundaries, which are typically encountered in the driving application of water desalination. We also stress that the coupling with salt transport adds complexity to the model. The unique solvability of the coupled problem is analysed by casting it as a fixed-point equation and using Banach’s fixed-point theorem. The fixed-point operator consists of the solution operator associated with the Navier–Stokes equations with membrane (or mixed slip-type) boundary conditions, composed with the solution operator associated with an advection-diffusion equation. The unique solvability of the first sub-problem is established using a Stokes linearisation combined with the fixed-point theory in the case of non-homogeneous boundary conditions. The unique solvability of the outer fixed-point problem results as a consequence of an assumption of smallness of data, which in our context reduces to impose a condition on the inlet velocity and on the constitutive equation for the membrane interaction term.
@@
We mention that related works where non-homogeneous mixed boundary conditions are of importance include, for instance, the solvability analysis of Navier–Stokes equations with free boundary [31], Boussinesq-type of equations with leaking boundary conditions and Tresca slip [28], the fixed-point analysis for Navier–Stokes equations with mixed boundary conditions [35], the work [24] addressing the analysis on viscous flows around obstacles with non-homogeneous boundary data, the solvability of Boussinesq equations with mixed (and non-homogeneous) boundary data [2, 11], and the regularity of split between normal and tangential parts of the velocity as boundary conditions for the Navier–Stokes equations in [14]. However it is important to mention that, up to our knowledge, the set of equations we face here has not been analysed in the existing literature.
The divergence-conforming discontinuous Galerkin (DG) method (introduced in [12]) represents a very useful numerical approach for solving partial differential equations, particularly in the context of fluid dynamics and electromagnetism. In contrast with the conforming formulation for incompressible flows and standard DG methods, the divergence-conforming variant ensures that the discrete velocity is divergence-free, a critical feature in view of conservative discretisations. Additionally, velocity error estimates could be determined in a manner that is resilient to variations in pressure. Moreover, locally, conservation guarantees a divergence-form representation for the coupled systems at the discrete level. Studies on this regard can be found in, for instance, [3, 37, 9, 38, 25]. In [37] the authors propose an div-free conforming scheme for a double diffusive flow on porous media, where the divergence-conforming Brezzi-Douglas-Marini (BDM) elements of order $k\geq 1$ are used for velocity approximation, discontinuous elements of order $k-1$ are used for the pressure, and continuous elements of order $k$ for the temperature. To enforce $\mathrm{H}^{1}$-continuity of velocities, they resort to an interior penalty DG technique. We also have the work of [9], in which the authors study a divergence-conforming finite element method for the doubly-diffusive problem. It considers temperature-dependent viscosity and potential cross-diffusion terms while maintaining the coercivity of the diffusion operator. The numerical scheme is based on H(div)-conforming BDM elements for velocity of order $k$, discontinuous elements for pressure of order $k-1$, and Lagrangian finite elements for temperature and solute concentration of order $k$. This formulation ensures divergence-free velocity approximations.
1.2 Plan
@@
The rest of the paper is organised as follows. The remainder of this section lists useful notation to be used throughout the paper. Section 2 is devoted to the governing equations and the specific boundary conditions needed in a typical operation of a desalination unit. There we also derive a weak formulation and provide preliminary properties of the weak forms. In Section 3 we conduct the analysis of existence and uniqueness of weak solution to the coupled system. We state an abstract result and show that the Navier–Stokes equations with membrane boundary conditions adhere to that setting. This section also describes the fixed-point analysis. Section 4 contains the definition of a conforming discretisation, a stabilisation technique proposed in [42], and then we define a new H(div)-conforming method and state main properties of the modified discrete variational forms. The unique solvability of the discrete problem is studied in Section 5, while the derivation of a priori error estimates is presented in Section 6. Qualitative properties of the proposed formulations are explored in Section 7, and we also confirm numerically the convergence rates predicted by the theory.
1.3 Preliminaries and notation
@@
Let us introduce some notations that will be used throughout the paper. Let $\Omega$ be a polygonal bounded domain of $\mathbb{R}^{2}$ with Lipschitz boundary $\partial\Omega$. We employ standard simplified notation for Lebesgue spaces, Sobolev spaces and their respective norms. Given $s\geq 0$ and $p\in[1,\infty]$, we denote by $\mathrm{L}^{p}(\Omega)$ Lebesgue space endowed with the norm $\|\cdot\|_{\mathrm{L}^{p}(\Omega)}$, while $\mathrm{H}^{s}(\Omega)$ denotes a Hilbert space. Vectors spaces and vector-valued functions will be written in bold letters. For instance, for $s\geq 0$, we simply write $\boldsymbol{\mathrm{H}}^{s}(\Omega)$ instead of $[\mathrm{H}^{s}(\Omega)]^{2}$. If $s=0$, we use the convention $\mathrm{H}^{0}:=\mathrm{L}^{2}(\Omega)$ and $\boldsymbol{\mathrm{H}}^{0}:=\mathbf{L}^{2}(\Omega)$. For the sake of simplicity, the seminorms and norms in Hilbert spaces are denoted by $|\cdot|_{s,\Omega}$ and $\|\cdot\|_{s,\Omega}$, respectively. The unit outward normal at $\partial\Omega$ is denoted by $\boldsymbol{n}:=(n_{1},n_{2})$, whereas $\boldsymbol{t}$ will denote the corresponding unit tangential vector perpendicular to $\boldsymbol{n}$ on $\partial\Omega$. Also, $\mathbf{0}$ denotes a generic null vector. Let us also define for $s\geq 0$ the Hilbert space $\mathrm{H}(\mathop{\mathrm{div}}\nolimits,\Omega):=\{\boldsymbol{v}\in\boldsymbol{\mathrm{H}}(\Omega):\mathop{\mathrm{div}}\nolimits\boldsymbol{v}\in\mathrm{H}(\Omega)\}$ whose norm is given by $\|\boldsymbol{v}\|_{\mathop{\mathrm{div}}\nolimits,\Omega}^{2}:=\|\boldsymbol{v}\|_{0,\Omega}^{2}+\|\mathop{\mathrm{div}}\nolimits\boldsymbol{v}\|_{0,\Omega}^{2}$. We also recall that for a Hilbert space $H$ with inner product $(\cdot,\cdot)_{H}$, $\mathcal{R}_{H}$ denotes the Riesz operator $H\rightarrow H^{\prime}$ that to each $z$ associates the functional $f_{z}=\mathcal{R}_{H}z\in H^{\prime}$ defined as
@@
with $\langle\cdot,\cdot\rangle_{H^{\prime}\times H}$ denoting the duality pairing between $H^{\prime}$ and $H$. $R_{H}$ is one-to-one, and $\|\mathcal{R}_{H}\|_{\mathcal{L}(H,H^{\prime})}=\|\mathcal{R}_{H}^{-1}\|_{\mathcal{L}(H^{\prime},H)}=1$. Moreover, if $H$ is identified with $(H^{\prime})^{\prime}$, then $\mathcal{R}_{H}^{-1}=\mathcal{R}_{H^{\prime}}$. Throughout the rest of the paper we abridge into $X\lesssim Y$ the inequality $X\leq CY$ with positive constant $C>0$ independent of $h$. Similarly for $X\gtrsim Y$.
@@ 2 Model problem
Let us consider that the boundary of $\Omega$ is decomposed as $\partial\Omega=\Gamma_{\mathrm{in}}\cup\Gamma_{\mathrm{out}}\cup\Gamma_{\mathrm{wall}}\cup\Gamma_{\mathrm{m}}$. The sub-boundary $\Gamma_{\mathrm{in}}$ corresponds to the inflow, $\Gamma_{\mathrm{out}}$ is the outflow boundary, $\Gamma_{\mathrm{wall}}$ is a no-slip no penetration boundary, and $\Gamma_{\mathrm{m}}$ is the porous membrane boundary, see Figure 2.1 for the particular case when $\Omega=[0,L]\times[0,d]$. The fluid inside the channel is assumed to be Newtonian, incompressible, and composed only by water and salt. The density is taken to be $\rho_{0}$, whereas the viscosity is given as $\mu_{0}$. The solute diffusivity through the solvent is given by $D_{0}$. It is also assumed that the effect of pressure drop inside the channel due to viscous effects on the permeate flux (solution-diffusion equation) is negligible.
Figure 2.1: Cross-flow membrane filtration model.
The resulting model is a coupling between Navier–Stokes and convection–diffusion equations: {linenomath}
@@
$\nabla\cdot\boldsymbol{u}$ $=0\quad\text{ in }\Omega,$ (1)
@@
Here, $\boldsymbol{u}(\boldsymbol{x})$, $p(\boldsymbol{x})$, and $\theta(\boldsymbol{x})$ represent the fluid velocity, pressure and concentration profile, respectively. Along with (2) we have the following set of boundary conditions: {linenomath}
[…]
In (2a), a parabolic inflow is considered with a fixed salt concentration representing a fully developed sea water flow trough the channel. In (2b)–(2c) we have a zero salt flux and do nothing boundary conditions. A zero Dirichlet boundary condition is imposed for the velocity across the impermeable wall. A full salt rejection is considered in the walls and the membrane, which is represented by (2e). In (2g) we have the permeability condition, where the quantity $g(\theta)$ denotes the flow velocity at the membrane as a function of the concentration and it can be represented using the Darcy–Starling law. As usual in membrane filtration processes, there are several orders of magnitude difference between the inlet and permeate flow velocities. More precisely, relating (2a) and (2g) we have the following inequality:
$0\leq g(\theta)\ll|\boldsymbol{u}_{\mathrm{in}}|.$ (3)
@@
$-\int_{\Gamma_{\mathrm{in}}}\rho_{0}\boldsymbol{u}_{\mathrm{in}}\cdot\boldsymbol{n}=\int_{\Gamma_{\mathrm{m}}\cup\Gamma_{\mathrm{out}}}\rho_{0}\boldsymbol{u}\cdot\boldsymbol{n}\qquad\text{and}\qquad 0\leq\int_{\Gamma_{\mathrm{m}}}g(\theta)\leq\int_{\Gamma_{\mathrm{out}}}\boldsymbol{u}\cdot\boldsymbol{n},$ (4)

### v2

@@ Abstract
We propose a new model for the coupling of flow and transport equations with porous membrane-type conditions on part of the boundary. The governing equations consist of the Navier–Stokes equations coupled with an advection-diffusion equation, and we employ a Lagrange multiplier to handle the coupling between penetration velocity and transport on the membrane, while mixed boundary conditions are considered in the remainder of the boundary. We show existence and uniqueness of the continuous problem using a fixed-point argument. Next, an H(div)-conforming finite element formulation is proposed, and we address its a priori error analysis. The method uses an upwind approach that provides stability in the convection-dominated regime. We showcase a set of numerical examples validating the theory and illustrating the use of the new methods in the simulation of reverse osmosis processes.
@@
Flow-transport equations; Lagrange multipliers; Reverse osmosis; Divergence-conforming FEM.
@@ 1 Introduction
The coupling of Navier–Stokes and advection-diffusion equations is central in many applications in industry and engineering. One of such instances is the simulation of filtration occurring in water purification, where high velocity flow with relatively high concentration of salt goes through a unit under reverse osmosis. Even without the mechanisms of ion transport and molecule interactions through the membrane, the fluid dynamics process already poses interesting challenges. In this simplified context, we can consider water and salt transport coupled through both advection and boundary interaction at the membrane – stating that velocity is proportional to the salt concentration. A similar model has been studied numerically in [9], where the coupling on the membrane follows Nitsche’s approach. Here we rewrite that model using a membrane Lagrange multiplier and provide a complete well-posedness analysis as well as the analysis of a divergence-conforming discretisation.
@@
Separating for a moment the salt transport from the incompressible flow equations, we end up with a generalisation of the Navier–Stokes equations with slip boundary conditions which have been studied extensively starting from the conforming finite element methods (FEMs) proposed in [38, 39]. The off-diagonal bilinear form is different from the usual one in that it also includes a pairing between the normal trace of velocity and the Lagrange multiplier taking into account the membrane coupling. Another difficulty in the model considered herein is that a non-homogeneous boundary condition is needed at the inlet. Often the analysis is simply restricted to the case of homogeneous essential boundary conditions, but here the non-homogeneity is important as it permits that the coupling occurs (otherwise the membrane coupling vanishes, and the only solution is the trivial one). We note that for smooth domains it is expected that discretisations are prone to the so-called Babuška paradox – a variational crime associated with the approximation of the boundary, and where a sub-optimal convergence is expected (see more details in, e.g., [23, 37]). Similar works focusing on slip boundary conditions imposed with Lagrange multipliers or with penalty can be found in [29, 27, 40]. In our case we restrict to polygonal boundaries, which are typically encountered in the driving application of water desalination.
We also stress that the coupling with salt transport adds complexity to the model. The unique solvability of the coupled problem is analysed by casting it as a fixed-point equation and using Banach’s fixed-point theorem. The fixed-point operator consists of the solution operator of the Navier–Stokes equations with membrane (or mixed slip-type) boundary conditions, composed with the solution operator associated with an advection-diffusion equation. The unique solvability of the first sub-problem is established using a Stokes linearisation combined with the fixed-point theory in the case of non-homogeneous boundary conditions. The unique solvability of the outer fixed-point problem results as a consequence of an assumption of smallness of data, which in our context reduces to impose a condition on the inlet velocity and on the constitutive equation for the membrane interaction term. Some parts of these proofs are standard but as we have not found them in the precise context we need, they have been included in the paper for sake of completeness.
@@
Related works where non-homogeneous mixed boundary conditions are of importance include, for instance, the solvability of Navier–Stokes equations with free boundary [28], Boussinesq equations with leaking boundary conditions and Tresca slip [25], the fixed-point analysis for Navier–Stokes equations with mixed boundary conditions [33], the analysis on viscous flows around obstacles with non-homogeneous boundary data [21], the solvability of Boussinesq equations with mixed (and non-homogeneous) boundary data [2, 10], and the regularity of split between normal and tangential parts of the velocity as boundary conditions for the Navier–Stokes equations in [14]. However it is important to mention that, up to our knowledge, the set of equations we face here has not been analysed in the existing literature.
The divergence-conforming discontinuous Galerkin (DG) method (introduced in [11]) represents a very useful numerical approach that produces divergence-free velocities. Additionally, velocity error estimates could be determined in a manner that is resilient to variations in pressure. Moreover, locally, conservation guarantees a divergence-form representation for the coupled systems at the discrete level. Studies on this regard can be found in, for instance, [3, 8, 35, 22].
Plan
@@
The rest of the paper is organised as follows. The remainder of this section lists useful notation to be used throughout the paper. Section 2 is devoted to the governing equations and the specific boundary conditions needed in a typical operation of a desalination unit. There we also derive a weak formulation and provide preliminary properties of the weak forms. In Section 3 we conduct the analysis of existence and uniqueness of weak solution to the coupled system. We state an abstract result and show that the Navier–Stokes equations with membrane boundary conditions adhere to that setting. This section also describes the fixed-point analysis. Section 4 contains the definition of a conforming discretisation, a stabilisation technique proposed in [39], and then we define a new H(div)-conforming method and state main properties of the modified discrete variational forms. This is accompanied by a brief discussion on a least-squares stabilised method using Lagrange multipliers, similar to that of [38, 39]. The unique solvability of the discrete divergence-free problem is studied in Section 5, while the derivation of a priori error estimates is presented in Section 6. Section 7 provides a brief description of a least-squares stabilised scheme using continuous elements. Qualitative properties of the proposed formulations are explored in Section 8, and we also confirm numerically the convergence rates predicted by the theory.
Preliminaries and notation
@@
Let $\Omega$ be a polygonal bounded domain of $\mathbb{R}^{2}$ with Lipschitz boundary $\partial\Omega$. We employ standard notation for Lebesgue spaces, Sobolev spaces and their respective norms. Given $s\geq 0$ and $p\in[1,\infty]$, we denote by $\mathrm{L}^{p}(\Omega)$ Lebesgue space endowed with the norm $\|\cdot\|_{\mathrm{L}^{p}(\Omega)}$, while $\mathrm{H}^{s}(\Omega)$ denotes a Hilbert space. Vectors spaces and vector-valued functions are written in bold letters. For instance, for $s\geq 0$, we simply write $\boldsymbol{\mathrm{H}}^{s}(\Omega)$ instead of $\mathrm{H}^{s}(\Omega)]^{2}$. If $s=0$, we use the convention $\mathrm{H}^{0}:=\mathrm{L}^{2}(\Omega)$ and $\boldsymbol{\mathrm{H}}^{0}:=\mathbf{L}^{2}(\Omega)$. For the sake of simplicity, the seminorms and norms in Hilbert spaces are denoted by $|\cdot|_{s,\Omega}$ and $\|\cdot\|_{s,\Omega}$, respectively. The unit outward normal at $\partial\Omega$ is denoted by $\boldsymbol{n}:=(n_{1},n_{2})$, whereas $\boldsymbol{t}$ denotes the corresponding unit tangential vector perpendicular to $\boldsymbol{n}$ on $\partial\Omega$. Also, $\mathbf{0}$ denotes a generic null vector. Let us also define for $s\geq 0$ the Hilbert space $\mathrm{H}(\mathop{\mathrm{div}}\nolimits,\Omega):=\{\boldsymbol{v}\in\mathbf{L}^{2}(\Omega):\mathop{\mathrm{div}}\nolimits\boldsymbol{v}\in\mathrm{L}^{2}(\Omega)\}$ whose norm is given by $\|\boldsymbol{v}\|_{\mathop{\mathrm{div}}\nolimits,\Omega}^{2}:=\|\boldsymbol{v}\|_{0,\Omega}^{2}+\|\mathop{\mathrm{div}}\nolimits\boldsymbol{v}\|_{0,\Omega}^{2}$. We also recall that for a Hilbert space $H$ with inner product $(\cdot,\cdot)_{H}$, $\mathcal{R}_{H}$ denotes the Riesz operator $H\rightarrow H^{\prime}$ that to each $z$ associates the functional $f_{z}=\mathcal{R}_{H}z\in H^{\prime}$ defined as
@@
with $\langle\cdot,\cdot\rangle_{H^{\prime}\times H}$ denoting the duality pairing between $H^{\prime}$ and $H$. $\mathcal{R}_{H}$ is one-to-one, and $\|\mathcal{R}_{H}\|_{\mathcal{L}(H,H^{\prime})}=\|\mathcal{R}_{H}^{-1}\|_{\mathcal{L}(H^{\prime},H)}=1$. Moreover, if $H$ is identified with $(H^{\prime})^{\prime}$, then $\mathcal{R}_{H}^{-1}=\mathcal{R}_{H^{\prime}}$. Throughout the rest of the paper we abridge into $X\lesssim Y$ the inequality $X\leq CY$ with positive constant $C>0$ independent of $h$. Similarly for $X\gtrsim Y$.
@@ 2 Model problem
Let us consider that the boundary of $\Omega$ is decomposed as $\partial\Omega=\Gamma_{\mathrm{in}}\cup\Gamma_{\mathrm{out}}\cup\Gamma_{\mathrm{wall}}\cup\Gamma_{\mathrm{m}}$. The sub-boundary $\Gamma_{\mathrm{in}}$ corresponds to the inflow, $\Gamma_{\mathrm{out}}$ is the outflow boundary, $\Gamma_{\mathrm{wall}}$ is a no-slip no penetration boundary, and $\Gamma_{\mathrm{m}}$ is the porous membrane boundary. The fluid inside the channel is assumed to be Newtonian, incompressible, and composed only by water and salt. The density and viscosity are constant and positive $\rho_{0},\mu_{0}>0$. The solute diffusivity through the solvent is given by $D_{0}$. It is also assumed that the effect of pressure drop inside the channel due to viscous effects on the permeate flux (solution-diffusion equation) is negligible.
Figure 2.1: Cross-flow membrane filtration model with $\Gamma_{i}$ defined such that there is no corner points on $\Gamma_{\mathrm{m}}\cap\Gamma_{\mathrm{in}}$.
The resulting model is a coupling between Navier–Stokes and convection–diffusion equations:
@@
(1) $\nabla\cdot\boldsymbol{u}$ $=0\quad\text{ in }\Omega,$
@@
Here, $\boldsymbol{u}(\boldsymbol{x})$, $p(\boldsymbol{x})$, and $\theta(\boldsymbol{x})$ represent the fluid velocity, pressure and concentration profile, respectively. Along with (2) we have the following set of boundary conditions:
(2a) $\boldsymbol{u}$ $=\boldsymbol{u}_{\mathrm{in}},\;\theta=\theta_{\mathrm{in}}$ $\text{ on }\Gamma_{\mathrm{in}},$
(2b) $-D_{0}\nabla\theta\cdot\boldsymbol{n}$ $=0$ $\text{ on }\Gamma_{\mathrm{out}},$
(2c) $\mu_{0}(\nabla\boldsymbol{u})\boldsymbol{n}-p\boldsymbol{n}$ $=\boldsymbol{0}$ $\text{ on }\Gamma_{\mathrm{out}},$
(2d) $\boldsymbol{u}$ $=\boldsymbol{0}$ $\text{ on }\Gamma_{\mathrm{wall}},$
(2e) $\left(\theta\boldsymbol{u}-D_{0}\nabla\theta\right)\cdot\boldsymbol{n}$ $=0$ $\text{ on }\Gamma_{\mathrm{m}}\cup\Gamma_{\mathrm{wall}},$
(2f) $\boldsymbol{u}\cdot\boldsymbol{t}$ $=0$ $\text{ on }\Gamma_{\mathrm{m}},$
(2g) $\boldsymbol{u}\cdot\boldsymbol{n}$ $=g(\theta)$ $\text{ on }\Gamma_{\mathrm{m}}.$
In (2a), a parabolic inflow is considered with a fixed salt concentration representing a fully developed sea water flow trough the channel. In (2b)–(2c) we have a zero salt flux and do nothing boundary conditions. A zero Dirichlet boundary condition is imposed for the velocity across the impermeable wall. A full salt rejection is considered in the walls and the membrane, which is represented by (2e). In (2g) we have the permeability condition, where the quantity $g(\theta)$ denotes the flow velocity at the membrane as a function of the concentration and it can be represented using the Darcy–Starling law. In order to enforce compatibility of the inlet and membrane boundary conditions we make precise the form of the parabolic profile as follows (assuming that the bottom left corner of the domain is located at the origin)
$\boldsymbol{u}_{\mathrm{in}}=\frac{1}{d}\bigl(u_{0}[d-y]y,-g(\theta_{\mathrm{in}})[d-y]\bigr)^{\tt t},\quad\boldsymbol{u}|_{\Gamma_{\mathrm{m}}}=(0,-g(\theta))^{\tt t},$
and we note that both $x$ and $y$ components of the velocity datum are continuous at the corner point $\Gamma_{\mathrm{in}}\cap\Gamma_{\mathrm{m}}$. This, or actually any smooth vertical inlet velocity that goes from $u_{\mathrm{in,2}}=-g(\theta)$ at the bottom left to $u_{\mathrm{in,2}}=0$ at the top left corner (and faster than linear) is a suitable choice for the analysis, which requires the velocity datum in $\boldsymbol{\mathrm{H}}^{1/2}(\Gamma_{\mathrm{in}}\cup\Gamma_{\mathrm{wall}}\cup\Gamma_{\mathrm{m}})$.
As usual in membrane filtration processes, there are several orders of magnitude difference between the inlet and permeate flow velocities. More precisely, relating (2a) and (2g) we have the following inequality:
(3) $0\leq g(\theta)\ll|\boldsymbol{u}_{\mathrm{in}}|.$
@@
(4) $\int_{\Gamma_{\mathrm{in}}}\rho_{0}\boldsymbol{u}_{\mathrm{in}}\cdot\boldsymbol{n}=\int_{\Gamma_{\mathrm{m}}\cup\Gamma_{\mathrm{out}}}\rho_{0}\boldsymbol{u}\cdot\boldsymbol{n}\qquad\text{and}\qquad 0\leq\int_{\Gamma_{\mathrm{m}}}g(\theta)\leq\int_{\Gamma_{\mathrm{out}}}\boldsymbol{u}\cdot\boldsymbol{n},$